OpenStudy (constanceleak):

PLS HELP!!! Will give medal! Two right triangles with the following dimensions are graphed with their hypotenuses on the same line: triangle X: horizontal leg measures 2 units vertical leg measures 8 units triangle Y: horizontal leg measures 4.5 units vertical leg measures 18 units Which other triangles could also have their hypotenuse on the same line?

OpenStudy (constanceleak):

Choose all answers that are correct. A. horizontal leg measures 3 units, vertical leg measures 12 units B. horizontal leg measures 5 units, vertical leg measures 18 units C. horizontal leg measures 4 units, vertical leg measures 16 units D. Horizontal leg measures 9 units, vertical leg measures 36 units

OpenStudy (kkutie7):

looks like you need to figure out the slope of the line on the graph.

OpenStudy (constanceleak):

Ok

OpenStudy (kkutie7):

Pythagorean theorem c^2=a^2+b^2. for both of the dimensions you have.

OpenStudy (constanceleak):

I'm in 8th grade... I don't think I know that yet.

OpenStudy (constanceleak):

i thought i was supposed to use y2-y1/x2-y1=m or slope

OpenStudy (kkutie7):

you can use equations like that to solve the slope yes, but just from what I know. is a tad bit easier to work with, but let me see.

OpenStudy (constanceleak):

ok. thx.

OpenStudy (constanceleak):

I think A,C, and D are right

OpenStudy (kkutie7):

 i imagine if you used your equation on both of them thinking of them this way you find a slope between the two.

OpenStudy (kkutie7):

using horizontal as x coordinates and vertical for y

OpenStudy (constanceleak):

ok. thx!

OpenStudy (kkutie7):

do you know that that is probably confusing. best bet is to graph them all on the same graph if you get stuck. goodlucky

OpenStudy (constanceleak):

ok.